Two rectangles have the same width. One is 12 units long and the other is 8 units long. The area of the first rectangle is 320 square units more than the area of the second rectangle. Find the width of each rectangle.
Answer by ankor@dixie-net.com(22740)

Two rectangles have the same width.
One is 12 units long and the other is 8 units long.
The area of the first rectangle is 320 square units more than the area of the second rectangle.
Find the width of each rectangle.
:
let w = the width of the rectangles
then
12w = the area of one rectangle
8w = the area of other
:
12w - 8w = 320
4w = 320
w = 320/4
w = 80 units wide
;
:
Check
12 * 80 = 960
8 * 80 = 640
--------------
difference: 320